OpenStudy (anonymous):

What two consecutive numbers have a sum of 53?

OpenStudy (shane_b):

\[n+(n+1)=53\]Just solve for n :)

OpenStudy (shane_b):

BTW: Obviously n will give you the first number...add +1 to get the second number.

OpenStudy (anonymous):

what's the first number?

OpenStudy (shane_b):

Did you solve this for n? \[n+(n+1)=53\]

OpenStudy (anonymous):

I don't remember how :(

OpenStudy (shane_b):

n+(n+1)=53 n+n+1=53 2n+1=53 2n+1-1=53-1 2n=52 52\2=n n=26